Hercules Exercise Equipment Co. purchased a computerized measuring device two years ago for $50,000. The equipment falls into the five-year category for MACRS depreciation and can currently be sold for $20,800. A new piece of equipment will cost $140,000. It also falls into the five-year category for MACRS depreciation. Assume the new equipment would provide the following stream of added cost savings for the next six years. Use Table 12–12. Use Appendix B for an approximate answer but calculate your final answer using the formula and financial calculator methods. Year Cash Savings 1 $ 58,000 2 50,000 3 48,000 4 46,000 5 43,000 6 32,000 The firm’s tax rate is 25 percent and the cost of capital is 8 percent. b. What is the tax loss on the sale of the old equipment? (Do not round intermediate calculations and round your answer to the nearest whole dollar.) c. What is the tax benefit from the sale? (Do not round intermediate calculations and round your answer to the nearest whole dollar.) d. What is the cash inflow from the sale of the old equipment? (Do not round intermediate calculations and round your answer to the nearest whole dollar.) e. What is the net cost of the new equipment? (Include the inflow from the sale of the old equipment.) (Do not round intermediate calculations and round your answer to the nearest whole dollar.) f. Determine the depreciation schedule for the new equipment. (Round the depreciation base and annual depreciation answers to the nearest whole dollar. Round the percentage depreciation factors to 3 decimal places.) g. Determine the depreciation schedule for the remaining years of the old equipment. (Round the depreciation base and annual depreciation answers to the nearest whole dollar. Round the percentage depreciation factors to 3 decimal places.) h. Determine the incremental depreciation between the old and new equipment and the related tax shield benefits. (Enter the tax rate as a decimal rounded to 2 decimal places. Round all other answers to the nearest whole dollar.) i. Compute the aftertax benefits of the cost savings. (Enter the aftertax factor as a decimal rounded to 2 decimal places. Round all other answers to the nearest whole dollar.) j-1. Add the depreciation tax shield benefits and the aftertax cost savings to determine the total annual benefits. (Do not round intermediate calculations and round your answers to the nearest whole dollar.) j-2. Compute the present value of the total annual benefits. (Do not round intermediate calculations and round your answer to the nearest whole dollar.) k-1. Compare the present value of the incremental benefits (j) to the net cost of the new equipment (e). (Do not round intermediate calculations. Negative amount should be indicated by a minus sign. Round your answer to the nearest whole dollar.)
Answer #1

Answer (b):

Hercules Exercise Equipment Co. purchased a computerized measuring device two years ago for $50,000.

Year 1 depreciation = 50000 * 20% = $10,000

Year 2 depreciation = 50000 * 32% =  $16,000

Accumulated depreciation after 2 years = 10000 + 16000 = $26,000

Book value now = 50000 - 26000 = $24,000

Machine can be sold at = $20,800

Loss on sale = 24000 - 20800 = $3,200

As the machine is sold incurring loss, tax loss on the sale of the old equipment = $0

Answer (c):

Tax benefit from the sale = 3200 * 25% = $800

Tax benefit from the sale = $800

Answer (d):

Cash inflow from the sale of the old equipment = Sale value + Tax benefit = 20800 + 800 = $21,600

Cash inflow from the sale of the old equipment = $21,600

Answer (e):

Net cost of the new equipment = Cost of new equipment - Sale proceeds from old equipment including tax benefit

= $140,000 - 21600

= $118,400

Net cost of the new equipment = $118,400

Answer (f):

5 6 New Machine Depreciation schedule: Year 1 Depreciation rate 20.000% Depreciation expense $ 28,000 $ 32.000% 44,800 $ 19.2

Answer (g):

Old Equipment Depreciation schedule for remaining life: Year (Old equipment life time line) Year (Timeline now) Depreciation

Answer (h):

6 8,064 Year Depreciation - New Equipment $ Depreciation - Old Equipment $ Incremental Depreciation $ Tax shield in increment

Answer (i):

Year Cash Savings Tax on cash savings at 25% Aftertax benefits of the cost savings. 1 $58,000 $14,500) $43,500 2 $50,000 ($12

Answer (j-1):

Year Aftertax benefits of the cost savings. Tax shield in incremental depreciation $ Total annual benefits. 1 $43,500 4,600$

Answer (j-2):

0 Year Aftertax benefits of the cost savings. Tax shield in incremental depreciation Total annual benefits. PV Factor at 8% (

Answer (k-1):

Present value of the total annual benefits Net cost of the new equipment Net Present Value $186,704.63 $118,400.00 $68,305

The project results in net present value of $68,305 and hence should be accepted.